OpenStudy (anonymous):

How do you express the following in algebraic expression: The value of a number increases by 22 every time another number doubles? I got: x+22/2y.

OpenStudy (dtan5457):

Why divide?

OpenStudy (anonymous):

was thinking in terms of X per y which's normally express as x/y example km/h. But I might be confusing my physics with math.

OpenStudy (anonymous):

The other one I have in mind is 2x = y+22 but it doesn't seem right.

OpenStudy (dtan5457):

This is harder than it looks,lol..

OpenStudy (anonymous):

I know right, I'm thinking it has some kind of factorial just can't remember the formula.

OpenStudy (anonymous):

How do express a number that doubles every time? Exponents right?

OpenStudy (dtan5457):

yes

OpenStudy (anonymous):

then how do you express a number increasing by 22 corresponding to the exponent of another number?

OpenStudy (dtan5457):

@Nnesha @Jhannybean

OpenStudy (dtan5457):

would there be more than 2 variables??

OpenStudy (anonymous):

I got a table here representing the growth but I've to express function... give me a moment to write it down.

OpenStudy (dtan5457):

alright. im really curious about this.

OpenStudy (anonymous):

1 = x; 2 = x+22; 4 = x+44; etc.

OpenStudy (anonymous):

That's it really, it asks for the table to be express as an algerbraicly i.e. x-110w.

OpenStudy (anonymous):

How about: \[x^{2} = y+22x\]

OpenStudy (dtan5457):

can't be..because that means that the number determines how many times you add 22..

OpenStudy (dtan5457):

if x=5 for example in that equation, it's doubled once but 22 times 5?

OpenStudy (anonymous):

You right, how about: \[x ^{y}=z+22y\]

OpenStudy (dtan5457):

the closest i can think of is \[x+22y=r^y \]

OpenStudy (anonymous):

lols, we got the same idea.

OpenStudy (dtan5457):

the problem with that is when it's doubled once.. if y=1 aka the number isn't doubled, the other side still adds the 22.

OpenStudy (dtan5457):

if it's doubled twice or y=2, then 22 gives it out 2 times instead of 1...

OpenStudy (anonymous):

\[If: x = 2, y \neq 1\]\[x ^{y} = z + 22y\]

OpenStudy (anonymous):

\[2^{5} = z +22(5)\]

OpenStudy (anonymous):

looks right to me...

OpenStudy (dtan5457):

How about... \[x+22(y-1)=r^y{-1}\]

OpenStudy (dtan5457):

double your equation once, replace y with 2

OpenStudy (dtan5457):

it doubles correctly on the exponent but gives 44 for only one double :/

OpenStudy (anonymous):

since 1 = x, 2 = x+22, 4 = x+44

OpenStudy (dtan5457):

my equation doesn't work either, lol.

OpenStudy (anonymous):

\[2^{2} = z + 22(2) = z\]

OpenStudy (anonymous):

z + 44

OpenStudy (dtan5457):

but the number doubled once

OpenStudy (dtan5457):

it's suppose to be z+22

OpenStudy (anonymous):

but 2 is already a double. The problem starts at 1. So: If x = 2, y /= 1 and x-1 = z then the equation stands, I think.

OpenStudy (anonymous):

Else we would need sigma

OpenStudy (dtan5457):

xy=z+22y 1 doesn't work ^2 is a double. y for 22 is suppose to =1, but it would be 2..

OpenStudy (dtan5457):

1 double=+22, not 44

OpenStudy (dtan5457):

Try this one

OpenStudy (dtan5457):

\[x+22(y-2)=r^{y-1}\]

OpenStudy (dtan5457):

That would work with a single double.

OpenStudy (dtan5457):

3-1=2 one double (3-2)=1 22

OpenStudy (anonymous):

lemme try it

OpenStudy (dtan5457):

it really looks like it works lol

OpenStudy (dtan5457):

As long as \[y \neq 0,1\]

OpenStudy (anonymous):

x is the base amount, y is the exponent and r is the number being doubled?

OpenStudy (anonymous):

just making sure I got the variables right...

OpenStudy (dtan5457):

x=the first number r=the 2nd number y to be honest does NOT represent the amount of times being doubled

OpenStudy (dtan5457):

the equation when being solved would put your statement true.

OpenStudy (dtan5457):

but y=any# that makes your equation true which is not 0,1

OpenStudy (dtan5457):

it doesn't really "stand" for anything

OpenStudy (anonymous):

r would colapse if it's 1

OpenStudy (dtan5457):

yes that's why y>1

OpenStudy (anonymous):

\[x+22(5-2)=1^{5-1}\]

OpenStudy (anonymous):

\[r \neq 1\]

OpenStudy (dtan5457):

r>1 y>1

OpenStudy (dtan5457):

to be honest this question is full of ****

OpenStudy (anonymous):

With that, my equation would have work. Lols, I love the challenge imo. Solving such pellet makes me feeling like Hawkins

OpenStudy (anonymous):

*problem

OpenStudy (dtan5457):

yes, but say your question doesn't allow restrictions then that question will end your life

OpenStudy (anonymous):

We can assume 1 = x, 2 = x + 22 since my equation starts at 2 then it's valid.

OpenStudy (anonymous):

We need 3rd person, we're just bouncing each others idea.

OpenStudy (dtan5457):

not sure what you mean by x=?

OpenStudy (anonymous):

\[2^{2} = 4; 2^{3} = 8; etc\]

OpenStudy (anonymous):

x = 1 is the base x + 22 = 2

OpenStudy (anonymous):

x + 44 = 4 x + 66 = 8

OpenStudy (anonymous):

hence: \[x + 22(y) = z ^{y}\]

OpenStudy (anonymous):

So what you think @JoannaBlackwelder

OpenStudy (joannablackwelder):

I'm working on it, just a sec.

OpenStudy (dtan5457):

your right the values would not equal to eachother

OpenStudy (joannablackwelder):

I don't think you can make this happen in an expression. It needs an equation.

OpenStudy (anonymous):

I think I got my vocabs wrong, I think I meant equations all along.

OpenStudy (anonymous):

The value of a number increases by 22 every time another number doubles: If \[z = 2, y \neq 1, z-1 = x\] then \[x + 22(y) = z ^{y}\]

OpenStudy (anonymous):

Just seems right to me but I'm not sure... \[1 = x\]\[2 = x + 22\]\[2^{2} = x+22(2) = 4 = x+44\]\[2^{3} = x+22(3) = 8 = x+66\]

OpenStudy (triciaal):